OpenStudy (anonymous):

what is x over 5=6.8

OpenStudy (anonymous):

\[\frac{ x }{ 5 } = 6.8\] \[x = 6.8*5\]

OpenStudy (anonymous):

every variable has a 1 behind it so it's 1x/5=6.8 multiply both sides by 5